on the hip, calculate how long it lasts. 2.27 BIO Are We Martians? It has been suggested, and not i facetiously, that life might have originated on Mars and been car- ried to the earth when a meteor hit Mars and blasted pieces of f rock (perhaps containing primitive life) free of the Martian sur- face. Astronomers know that many Martian rocks have come to the earth this way. (For instance, search the Internet for ALH 84001.) One objection to this idea is that microbes would have had to undergo an enormous lethal acceleration during the impact. Let us investigate how large such an acceleration might be. To escape Mars, rock fragments would have to reach its escape velocity of 5.0 km/s, and that would most likely happen over a distance of about 4.0 m during the meteor impact. (a) What would be the acceleration (in m/s2 and gs) of such a rock frag- ment, if the acceleration is constant? (b) How long would this acceleration last? (c) In tests, scientists have found that over 40% of Bacillus subtilis bacteria survived after an acceleration of 450,000g. In light of your answer to part (a), can we rule out the hypothesis that life might have been blasted from Mars to the earth? つつ。
Answer #1

A)

Acceleration, a =v^2/2d = (5000)^2/2 x 4 = 3125000 m/s^2

B)

Acceleration, a = 3125000/9.81 = 318552.5 g

C)

Time, t = v/a = (5000)/3125000 = 0.0016 sec = 1.6 ms

Comment in case any doubt please rate my answer ...